REU 2007: SCHEDULE FOR WEEKS FIVE AND SIX

AFTERNOON TALKS ARE IN ECKHART 206

FIFTH WEEK

Date YSP Bader Fiore Angeltveit
Mon, July 16 9:30 – 2:30 4:30 – 5:30 3:00 – 4:00
Tue, July 17 9:30 – 2:30  3:00 – 4:00 4:30 – 6:00
Wed, July 18 9:30 – 2:30 4:30 – 5:30 3:00 – 4:00
Thu, July 19 9:30 – 2:30  3:00 – 4:00 4:30 – 6:00
Fri, July 20 9:30 – 2:30 4:30 – 5:30 3:00 – 4:00

SIXTH WEEK

Date YSP Babai Bader Fiore Angeltveit
Mon, July 23 9:30 – 2:30 4:30 – 6:00 3:00 – 4:00
Tue, July 24 9:30 – 2:30  3:00 – 4:00 4:30 – 6:00
Wed, July 25 9:30 – 2:30 4:30 – 6:00 3:00 – 4:00
Thu, July 26 9:30 – 2:30  3:00 – 4:00 4:30 – 6:00
Fri, July 27 9:30 – 2:30 4:30 – 6:00 3:00 – 4:00

Here is the schedule in summary form:

FIFTH WEEK:

YSP 9:30 – 2:30 M-F Bader (Groups of matrices) 4:30 – 5:30 MWF Fiore (Mathematics and music) 3:00 – 4:00 M-F Angeltveit (Products on spaces) 4:30 – 6:00 TTh

SIXTH WEEK:

YSP 9:30 – 2:30 M–F Babai (Discrete) 4:30 – 6:00 MWF Bader (Groups of matrices) 3:00 – 4:00 MWF Fiore (Mathematics and music) 3:00 – 4:00 TTh Angeltveit (Products on spaces) 4:30 – 6:00 TTh